# TrajAgent: An LLM-based Agent Framework for Automated Trajectory Modeling via Collaboration of Large and Small Models

### v3

@@ 3.4. Comparison with Automated Methods
Here, we compare TrajAgent with the deep learning-based AutoML method Optuna (Akiba et al., 2019) and the LLM-based automated urban task-solving framework UrbanLLM (Jiang et al., 2024). Using the trajectory user linking task with the S2TUL model as an example, the results are presented in Figure 5. UrbanLLM (represented by the blue line) is designed to directly automate the use of existing models without further optimization, resulting in the lowest performance in Figure 5. Meanwhile, compared to the widely used deep learning-based AutoML method Optuna (represented by the red line), which focuses on parameter optimization, TrajAgent (depicted by the yellow line) achieves superior results with fewer trial-and-error iterations. Furthermore, its performance can be further enhanced through joint optimization combined with automated data augmentation, resulting in a significant performance improvement of over 11.1%.
4. Related Work
4.1. Trajectory Modelling and Analytics
@@
In recent year, trajectory modelling (Zheng, 2015; Chen et al., 2024; Graser et al., 2024) makes great progress on its core research questions, including prediction (Liu et al., 2016; Feng et al., 2018; Shen et al., 2020), classification (Liang et al., 2022; Song et al., 2018), recovery (Xia et al., 2021) and generation (Yuan et al., 2022). Trajectory prediction is to predict the next location or the trip destination of the trajectory. ST-RNN (Liu et al., 2016) is the first work for applying recurrent neural network into the trajectory prediction task. Following this work, lots of works have been done to solve the critical issues of trajectory prediction with deep learning, for example, DeepMove (Feng et al., 2018) for multi-granularity regularity modelling, Flash-back (Yang et al., 2020) for sparse trajectory prediction, GETNext (Yang et al., 2022) for utilizing aggregated mobility pattern. As for the classification problem in trajectory modelling, the most well-known tasks are trajectory mode detection (Jiang et al., 2017) and trajectory user linkage (Gao et al., 2017) task. Various techniques, such as siamese network (Feng et al., 2019), distillation learning (Chen et al., 2022) and transformers (Liang et al., 2022) are successfully applied to discover the pattern among trajectories to solve the classification tasks. Finally, due to the natural limitation of trajectory data quality, trajectory data recovery (Xi et al., 2019; Xia et al., 2021) and generation (Ouyang et al., 2018) are also important tasks in trajectory modelling which are crucial for the practical application in the real world. Thus, follow the the trend of generative modelling, diverse generative models including GAN (Ouyang et al., 2018), diffusion models (Yuan et al., 2022; Wei et al., 2024) are utilized to improve the modelling of human dynamics. While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task. In other words, the automated and unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data. In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various trajectory data and trajectory modelling tasks without human intervention.
4.2. Large Language Model and Agent
With the advent of ChatGPT (OpenAI, 2024), large language models(LLMs) with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics (Yao et al., 2022b), question answering (Zhuang et al., 2023), and human-machine interaction (Touvron et al., 2023). Furthermore, agentic framework (Wang et al., 2024; Sumers et al., 2023; Xi et al., 2023) are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024). Following this promising direction, researchers also explore the potential of applying LLM based agent for automatic research experiments (Bogin et al., 2024; Romera-Paredes et al., 2024; Boiko et al., 2023) especially machine learning experiments (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024b). For example, MLAgentBench (Huang et al., 2024) designs MLagent to automatically solve simple kaggle competition with ReAct mechanism in terms of success rate. DSBench (Zhang et al., 2024) verifies the LLM based agent for solving classic data science tasks with tabular data as input. However, trajectory data are more complicated than the tabular data or text data and different trajectory modelling tasks also exhibit diverse problem settings and models, which are too challenging for these existing agents designed for general machine learning tasks with simple data input and interface. Recently, UrbanLLM (Jiang et al., 2024) follows the idea of HuggingGPT (Shen et al., 2024) to manage the existing specialized spatial-temporal models and APIs for tackling urban related tasks, which covers some trajectory modelling tasks. However, existing specialized spatial-temporal AI models usually designed for specific task with specific data and format, directly utilizing these models can suffer significant performance drop even can not generate any make sense results. In this paper, different from general machine learning agent and UrbanLLM for static spatial-temporal model, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
5. Conclusion
@@
In this paper, we propose TrajAgent, an LLM-based agentic framework for automated trajectory modeling. Supported by UniEnv, which provides a unified data and model interface, and collaborative learning schema for joint performance optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modeling tasks. TrajAgent introduces a new paradigm for unified trajectory modeling across diverse datasets and tasks, paving the way for advancements in trajectory modeling applications.
@@ A.1. Datasets
  • Tencent (Liu et al., 2023): This dataset includes both a road network and a set of vehicle trajectories collected in northeastern Beijing. The road network consists of 8.5K road segments and 15K edges, and the trajectory dataset contains 64K vehicle trajectories, each sampled at 15-second intervals.
  • Beijing (Li et al., 2024): This dataset contains check-in records collected from a popular social networking platform. It spans a period from late September 2019 to late November 2019.It also contains intent labels annotated by human for a small dataset.
  • UserQueries: To verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
@@
Table 5. Statistics of the Check-in trajectory datasets.
@@
Table 6. Statistics of the GPS trajectory datasets.
@@ A.2. Models
  • Next Location Prediction: RNN (Liu et al., 2016), attention-based method like DeepMove (Feng et al., 2018), well-performed graph-based method like GETNext (Yang et al., 2022) and two recent proposed LLM-based methods LLM-ZS (Beneduce et al., 2024) and LLMMove (Feng et al., 2024).
  • Travel Time Estimation: DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a) to estimate the travel time for a given GPS trajectory.
  • Trajectory User Linkage: widely-used DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022) and S2TUL (Deng et al., 2023) are considered. We modified DPLink’s training approach by using publicly available sparse trajectory datasets instead of heterogeneous mobility datasets for training.
  • Travel Intent Prediction: LIMP (Li et al., 2024), which leverages the commonsense reasoning capabilities of LLMs for mobility intention inference.
  • Trajectory Anomaly Detection: we consider the well-performing method GMVSAE (Liu et al., 2020b), which represent different types of normal trajectories in a continuous latent space.
  • Trajectory Generation: ActSTD (Yuan et al., 2022), which capture the spatiotemporal dynamics underlying trajectories by leveraging neural differential equations and DSTPP (Yuan et al., 2023), which defines spatial temporal point process for trajectory generation.
  • Trajectory Recovery: TrajBERT (Si et al., 2023), which encode trajectory as sentence and train a BERT model to get representations of trajectories.
  • Trajectory Map Matching: DeepMM (Feng et al., 2020), which proposes a data augmentation approach for map-based trajectory data, and GraphMM (Liu et al., 2023), which leverages a graph-based framework to extract features from map-based trajectory data, are considered.
  • Trajectory Representation: CASCSR (Gong et al., 2023), which use contrastive learning method to learn trajectory representations for downstream tasks.
A.3. Metrics
@@
In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, it is also used to measure the downstream applications of trajectory representation.$Acc_{m}$ is designed for measuring the performance of map matching task.$Acc_{i}$ is designed for measuring the performance of travel intention prediction task. Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation and trajectory recovery, while the metric AUC is used to assess the performance of trajectory anomaly detection.JSD and RMSE are used to measure the prediction error of the spatiotemporal domain in trajectory generation task.
A.4. Additional results on GPS trajectory data
Table 7. Comparison of task performance on GPS trajectories for TrajAgent with different configurations.
@@ A.4. Additional results on GPS trajectory data
A.5. Case Study
@@
The first issue is optimization trap present in data augmentation module of TrajAgent. Specifically, it refers to the situation where agent sometimes ignores the contents of the memory during the thought process. Even when the chosen parameter combination yields poor training results, the model overlooks the error feedback (i.e., the ”Not good enough…” in the memory). The ”optimization trap” occurs even in the best-performing GPT-4o-mini. As the length of the Memory increases, the impact of the optimization trap on overall accuracy grows. Once an optimization trap occurs, all memories within the same step tend to favor the same ineffective combination. We believe the causes of the optimization trap could be: (1) excessively long prompts leading to truncated inputs; (2) insufficient proportion of memory in the total input.

### v5

@@ 3.5 Comparison with Automated Methods
We compare TrajAgent with the deep learning-based AutoML method Optuna akiba2019optuna and the LLM-based automated urban task-solving framework UrbanLLM jiang2024urbanllm . Using the trajectory user linking task with S2TUL model as an example, the results are presented in Figure LABEL:fig:optuna. UrbanLLM (represented by the blue line) is designed to directly automate the use of existing models without further optimization, resulting in the lowest performance in Figure LABEL:fig:optuna. Meanwhile, compared to the widely used AutoML method Optuna (represented by the red line), which focuses on parameter optimization, TrajAgent (depicted by the yellow line) achieves superior results with fewer trial-and-error iterations. Furthermore, its performance can be further enhanced through joint optimization combined with automated data augmentation, resulting in a significant performance improvement of over 11.1%. For computational overhead please of TrajAgent, please refer to Table 8 in Appendix.
4 Related Work
@@
Trajectory Modelling and Analytics: In recent year, trajectory modelling zheng2015trajectory ; chen2024deep ; graser2024mobilitydl makes great progress on its core research questions, including prediction liu2016predicting ; feng2018deepmove ; yang2020location ; shen2020ttpnet ; yang2022getnext , classification jiang2017trajectorynet ; gao2017identifying ; feng2019dplink ; liang2022trajformer ; song2018anomalous ; chen2022mutual , recovery xi2019modelling ; xia2021attnmove and generation ouyang2018non ; yuan2022activity ; wei2024diff . While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task. In other words, the automated and unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data. In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various data and modelling tasks without human intervention.
Large Language Models: LLMs with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics yao2022react , question answering zhuang2023toolqa , and human-machine interaction touvron2023llama . Following this direction and motivated by the exploration of LLMs’ usability in urban studies feng2025citybench , researchers have begun developing diverse domain-specific LLMs tailored for urban applications, for example, CityGPT feng2025citygpt , UrbanLLM jiang2024urbanllm , and UrbanLLaVA feng2025urbanllava . In contrast to these approaches, which rely on fine-tuning LLMs with domain-specific knowledge, our work focuses on a training-free paradigm by constructing an agentic framework.
LLM based Agents: In the general domain, agentic framework wang2024survey ; sumers2023cognitive ; xi2023rise are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation yao2022webshop ; nakano2021webgpt ; zhou2023webarena and software development hong2023metagpt ; qian2024chatdev ; yang2024swe . Besides, researchers also explore the potential of applying LLM based agent for automatic research experiments bogin2024super ; romera2024mathematical ; boiko2023autonomous especially machine learning experiments huang2024mlagentbench ; shen2024hugginggpt ; zhang2024benchmarking ; wu2024visionllm . Recently, LLM-based agent also be applied in specific trajectory modeling tasks, e.g., trajectory prediction feng2024agentmove and trajectory simulation du2025cams . In this paper, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
5 Conclusion
@@
In this paper, we propose TrajAgent, an LLM-based agentic framework for automated trajectory modeling. Supported by UniEnv, which provides a unified data and model interface, and collaborative learning schema for joint performance optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modeling tasks. TrajAgent establishes a new paradigm for unified trajectory modeling across diverse tasks and datasets.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant 2024YFC3307603, in part by the China Postdoctoral Science Foundation under grant 2024M761670 and GZB20240384, in part by the Tsinghua University Shuimu Scholar Program under grant 2023SM235. This research is supported by Tsinghua University-Mercedes Benz Institute for Sustainable Mobility.
@@ A.1 Datasets
  • Tencent [34]: This dataset includes both a road network and a set of vehicle trajectories collected in northeastern Beijing. The road network consists of 8.5K road segments and 15K edges, and the trajectory dataset contains 64K vehicle trajectories, each sampled at 15-second intervals.
  • Beijing [28]: This dataset contains check-in records collected from a popular social networking platform. It spans a period from late September 2019 to late November 2019.It also contains intent labels annotated by human for a small dataset.
  • UserQueries: To verify the effectiveness of the whole system, we utilize self-instruct method [50] with 5 seed queries to generate 300 user queries as the experiment input.
@@
Table 5: Statistics of the Check-in trajectory datasets.
@@
Table 6: Statistics of the GPS trajectory datasets.
@@ A.2 Models
  • Next Location Prediction: RNN [31], attention-based method like DeepMove [13], well-performed graph-based method like GETNext [60] and two recent proposed LLM-based methods LLM-ZS [2] and LLMMove [19].
  • Travel Time Estimation: DeepTTE [48] and MulT-TTE [32] to estimate the travel time for a given GPS trajectory.
  • Trajectory User Linkage: widely-used DPLink [18], MainTUL [5] and S2TUL [10] are considered. We modified DPLink’s training approach by using publicly available sparse trajectory datasets instead of heterogeneous mobility datasets for training.
  • Travel Intent Prediction: LIMP [28], which leverages the commonsense reasoning capabilities of LLMs for mobility intention inference.
  • Trajectory Anomaly Detection: we consider the well-performing method GMVSAE [33], which represent different types of normal trajectories in a continuous latent space.
  • Trajectory Generation: ActSTD [65], which capture the spatiotemporal dynamics underlying trajectories by leveraging neural differential equations and DSTPP [64], which defines spatial temporal point process for trajectory generation.
  • Trajectory Recovery: TrajBERT [44], which encode trajectory as sentence and train a BERT model to get representations of trajectories.
  • Trajectory Map Matching: DeepMM [14], which proposes a data augmentation approach for map-based trajectory data, and GraphMM [34], which leverages a graph-based framework to extract features from map-based trajectory data, are considered.
  • Trajectory Representation: CASCSR [21], which use contrastive learning method to learn trajectory representations for downstream tasks.
A.3 Metrics
@@
[…]
All experiments are conducted on a Ubuntu server equipped with 8 NVIDIA GeForce RTX 3090 GPUs. Each small model is trained using a single RTX 3090 GPU, while each large language model (LLM) is accessed through its corresponding API provider.
A.4 Additional results on GPS trajectory data
Table 7: Comparison of task performance on GPS trajectories for TrajAgent with different configurations.
@@ A.4 Additional results on GPS trajectory data
A.5 Limitations and Failure Mode Analysis
@@
The first issue is optimization trap present in data augmentation module of TrajAgent. Specifically, it refers to the situation where agent sometimes ignores the contents of the memory during the thought process. Even when the chosen parameter combination yields poor training results, the model overlooks the error feedback (i.e., the "Not good enough…" in the memory). The "optimization trap" occurs even in the best-performing GPT-4o-mini. As the length of the Memory increases, the impact of the optimization trap on overall accuracy grows. Once an optimization trap occurs, all memories within the same step tend to favor the same ineffective combination. We believe the causes of the optimization trap could be: (1) excessively long prompts leading to truncated inputs; (2) insufficient proportion of memory in the total input.